# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ 3 Vector valued extensions of operators on convex body valued functions
$|W(x)^{1/p}G(x)|\sim_{d,p}\left(\sum_{n=1}^{\infty}|W(x)^{1/p}G_{n}(x)|^{q}\right)^{1/q}$
at almost every $x\in\mathbb{R}^{n},$ for any $1<p<\infty.$ It is also easy to see that the convex body valued functions $F,G$ are measurable.
@@
Here we have used Theorem A with the extrapolation paris $(F_{n},G_{n})\in\mathcal{F}$ and with exponent $q,$ so that we have
@@ 3.1 Fefferman–Stein vector valued inequalities for weighted maximal operators
where $f$ is a locally integrable function taking values on $\mathbb{C}^{d}$ and where the supremum is taken over all cubes with sides parallel to the coordinate axes. The maximal operator $M_{W}$ is also called Christ–Goldberg maximal operator, since it was studied by these two authors (see [Christ_Goldberg_2001] and [Goldberg_2003]). One can also define a weighted maximal operator with reducing operators as
@@
Both weighted maximal operators play an important role in the theory of matrix weighted norm inequalities. In addition, whenever $W$ is a matrix $A_{p}$ weight, both maximal operators are bounded from weighted $\mathrm{L}^{p}(W)$ to unweighted $\mathrm{L}^{p}.$ Isralowitz and Moen [Isralowitz_Moen_2019] proved that
@@
while Isralowitz, Kwon and Pott [Isralowitz_Kwon_Pott2017] showed that
@@
We will use a trick due to Bownik and Cruz-Uribe [Cruz_Uribe_Bownik_Extrapolation] so as to use Theorem 5 (which is a statement about convex body valued functions) to show statements about $\mathbb{C}^{d}$ vector valued functions. In order to apply the previous results for convex body valued operators to these weighted maximal operators, we will also need to define the convex body valued analogue of $M_{W}.$ Both definitions will be related by the following correspondence between vector valued and convex body valued functions. Given a vector valued function $f,$ that is taking values on $\mathbb{C}^{d},$ we define the convex body valued function $F$ by
@@
where the union is taken over all bounded cubes with sides parallel to the axes. Moreover, given a locally integrable vector valued function $f,$ we denote $M_{W}^{\mathcal{K}}f(x)=M_{W}^{\mathcal{K}}F(x),$ where $F$ is the convex body valued function given by (3.6). In addition, if $W$ is a matrix weight and $f$ a vector valued function, linearity and (3.8) yield that
@@
for any $1<p,q_{1}<\infty$ (with $q_{1}$ possibly different of $q=p_{0}$) and any matrix weight $W\in A_{p},$ where $\{(M^{\mathcal{K}}F_{n},F_{n})\}$ is a sequence in $\mathcal{F}$ and where
@@
Finally, replacing (3.11) and (3.12) into (3.10), we get the desired vector valued estimates for the pointwise weighted maximal operator, as we wanted to show. ∎
The analogous result for the operator $\widetilde{M}_{W}$ is a consequence of Theorem 1.
@@
An application of the classical Fefferman–Stein vector valued inequalities for the maximal operator followed by the use of Proposition 1 yields the desired result. ∎
@@ 4 Two matrix weighted biparameter product BMO
Let $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ be any product dyadic grid in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$
Let $B=\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ be any sequence in $\mathrm{M}_{d}(\mathbb{C}).$ We emphasize that $\mathcal{E}$ stands for the set of biparameter signatures, $\mathcal{E}=\mathcal{E}^{1}\times\mathcal{E}^{2},$ where $\mathcal{E}^{1}=\{0,1\}^{n}\setminus\{(1,\ldots,1)\}$ and $\mathcal{E}^{2}=\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$ We define
@@
$|\mathcal{V}_{R}P\mathcal{U}^{-1}_{R}|=|\mathcal{U}^{-1}_{R}P^{\ast}\mathcal{V}_{R}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P^{\ast}(\mathcal{V}^{\prime}_{R})^{-1}|,\qquad\forall R\in\bm{\mathcal{D}},\leavevmode\nobreak\ \forall P\in\mathrm{M}_{d}(\mathbb{C}).$
@@ 4.1 $\mathrm{H}^{1}$-BMO duality
The main goal of this subsection is to prove Theorem 2. We split the proof in Propositions 8 and 9 below, each treating one direction of Theorem 2.
@@
This is the direct biparameter analog of the one-parameter two matrix weighted $\mathrm{H}^{1}$ norm defined in [Isralowitz_2017]. It is easy to check that $(\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),\|\,\cdot\,\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)})$ is a Banach space.
@@
Let $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p).$ Then, the linear functional $\ell_{B}:\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)\rightarrow\mathbb{C},$
@@
is well-defined bounded with $\|\ell_{B}\|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}.$
@@
We adapt the first half of the proof of [Isralowitz_2017, Theorem 1.3]. Let $\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ be arbitrary. We show that the sum
$(\Phi,B)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})$
@@
$|(\Phi,B)|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}.$
@@
$\lesssim_{p,d}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|.$
Set now
@@
$\mathcal{B}_{k}\coloneq\left\{R\in\bm{\mathcal{D}}:\leavevmode\nobreak\ |R\cap\Omega_{k+1}|\leq\frac{1}{2}|R|<|R\cap\Omega_{k}|\right\},\qquad k\in\mathbb{Z},$
@@
Clearly $\Omega_{k}\subseteq\widetilde{\Omega}_{k}$ up to a set of zero measure, and in fact $|\widetilde{\Omega}_{k}|\sim|\Omega_{k}|,$ for all $k\in\mathbb{Z}.$ It is also obvious that $R\subseteq\widetilde{\Omega_{k}},$ for all $R\in\mathcal{B}_{k},$ for all $k\in\mathbb{Z}.$ If $R\in\bm{\mathcal{D}}$ with $\Phi_{R}^{\varepsilon}\neq 0$ for some $\varepsilon\in\mathcal{E}$ and $|R|\geq 2|R\cap\Omega_{k}|,$ for all $k\in\mathbb{Z},$ then by letting $k\rightarrow-\infty$ we deduce
$|R|\geq 2|R\cap\{F>0\}|=2|R|,$
since $F>0$ a. e. on $R,$ contradiction. If $R\in\bm{\mathcal{D}}$ with $|R|<2|R\cap\Omega_{k}|,$ for all $k\in\mathbb{Z},$ then by letting $k\rightarrow\infty$ we deduce
$|R|\leq 2|R\cap\{F=\infty\}|=0,$
since by assumption $F\in\mathrm{L}^{1}(\mathbb{R}^{n+m}),$ contradiction. It follows that for every $R\in\bm{\mathcal{D}}$ with $\Phi_{R}\neq 0$ for some $\varepsilon\in\mathcal{E}$ there exists $k\in\mathbb{Z}$ such that $R\in\mathcal{B}_{k}.$ Thus, we have
$|(\Phi,B)|$ $\lesssim_{p,d}\sum_{k\in\mathbb{Z}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|$
@@
This will be enough to conclude the proof, because assuming it we will get
$\sum_{k\in\mathbb{Z}}\bigg{(}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\bigg{)}^{1/2}|\widetilde{\Omega}_{k}|^{1/2}\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\widetilde{\Omega}_{k}|$
$\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\Omega_{k}|\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|F\|_{\mathrm{L}^{1}(\mathbb{R}^{n+m})}.$
@@
by the definition of $\Omega_{k+1}.$ Moreover, we have
$\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}F(x)^{2}\,\mathrm{dm}(x)\geq\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\,\mathrm{dm}(x)$
$=\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\,\mathrm{dm}(x)$
$\sim_{d}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{2}\,\mathrm{dm}(x)$
$\geq\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{|R\setminus\Omega_{k+1}|}{|R|}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}.$
@@
$[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}\lesssim_{p,d}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{p^{\prime}/p},$
because $V^{-1/p}$ is a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight (see for example Lemma 3.2 in [DKPS2024]). It is then well-known that
$[w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}\leq[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{1/p^{\prime}}$
@@
(see for example Subsection 2.3.3 in [Kakaroumpas_Soler_2022]). Using Jensen’s inequality and the definition of the $A_{2}$ characteristic, it follows that
@@
Thus
$\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$
$\gtrsim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{-2/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$

### v3

@@ 3 Vector valued extensions of operators on convex body valued functions
$|W(x)^{1/p}G(x)|\sim_{d,p}\left(\sum_{m=1}^{\infty}|W(x)^{1/p}G_{m}(x)|^{q}\right)^{1/q}$
at almost every $x\in\mathbb{R}^{n},$ for any $1<p<\infty.$
Let us prove that the convex body valued functions $F,G$ are measurable. For each $m=1,2,\ldots$ define the convex body valued function $F^{m}:\mathbb{R}^{n}\to\mathcal{K}_{\mathrm{bcs}}$ by
$F^{m}(x):=\Sigma_{q}(\{F_{t}(x)\}^{m}_{t=1}),\quad x\in\mathbb{R}^{n}.$
Fix $m$. For each $t=1,\ldots,m$, by using the characterization (2.6) of measurability of convex body valued functions we can find a sequence $\{f_{k}^{t}\}_{k=1}^{\infty}$ of measurable selections functions of $F_{t}$ such that
$F_{t}(x)=\overline{\{f_{k}^{t}(x):\penalty 10000\ k=1,2,\ldots\}}.$
Then, it is easy to see that
$F^{m}(x)=\overline{\left\{\sum_{t=1}^{m}a_{t}f^{t}_{k_{t}}(x)\colon k_{t}\in\mathbb{N},\penalty 10000\ a_{t}\in\mathbb{Q}\text{ for }t=1,\ldots,m\text{ with }\sum_{t=1}^{m}|a_{t}|^{q^{\prime}}\leq 1\right\}}.$
Thus, by the characterization (2.6) of measurability of convex body valued functions coupled with the fact that $\mathbb{N}^{m}$ is still a countable set, it follows that each $F^{m}$ is measurable.
Moreover, it is clear that
$F(x)=\overline{\bigcup_{m=1}^{\infty}F_{m}(x)}$
and therefore also
$F(x)=\overline{\mathrm{conv}\left(\bigcup_{m=1}^{\infty}F_{m}(x)\right)},$
since $F(x)$ is already known to be a closed convex set. Thus by [1, Theorem 8.2.2] we deduce that $F$ is measurable. Similarly, $G$ is also measurable.
@@
Here we have used Theorem A with the extrapolation pairs $(F_{n},G_{n})\in\mathcal{F}$ and with exponent $q,$ so that we have
@@ 3.1 Fefferman–Stein vector valued inequalities for weighted maximal operators
where $f$ is a locally integrable function taking values on $\mathbb{C}^{d}$ and where the supremum is taken over all cubes with sides parallel to the coordinate axes. The maximal operator $M_{W}$ is called Christ–Goldberg maximal operator, since it was introduced and studied by these two authors (see [7] and [20]). One can also define a weighted maximal operator with reducing operators as
@@
Both weighted maximal operators play an important role in the theory of matrix weighted norm inequalities. In addition, whenever $W$ is a matrix $A_{p}$ weight, both maximal operators are bounded from weighted $\mathrm{L}^{p}(W)$ to unweighted $\mathrm{L}^{p}.$ Isralowitz and Moen [26] proved that
@@
while Isralowitz, Kwon and Pott [25] showed that
@@
We will use a trick due to Bownik and Cruz-Uribe [5] so as to use Theorem 4 (which is a statement about convex body valued functions) to show statements about $\mathbb{C}^{d}$ vector valued functions. In order to apply the previous results for convex body valued operators to these weighted maximal operators, we will also need to define the convex body valued analogue of $M_{W}.$ Both definitions will be related by the following correspondence between vector valued and convex body valued functions. Given a vector valued function $f,$ that is taking values on $\mathbb{C}^{d},$ we define the convex body valued function $F$ by
@@
where the union is taken over all bounded cubes with sides parallel to the axes. Moreover, given a locally integrable vector valued function $f,$ we denote $M_{W}^{\mathcal{K}}f(x)=M_{W}^{\mathcal{K}}F(x),$ where $F$ is the convex body valued function given by (3.8). In addition, as proved in [41, Lemma 3.4], if $W$ is a matrix weight and $f$ a vector valued function, linearity and (3.10) yield that
@@
for any $1<p,q_{1}<\infty$ (with $q_{1}$ possibly different if $q=p_{0}$) and any matrix weight $W\in A_{p},$ where $\{(M^{\mathcal{K}}F_{n},F_{n})\}$ is a sequence in $\mathcal{F}$ and where
@@
Finally, replacing (3.13) and (3.14) into (3.12), we get the desired vector valued estimates for the Christ–Goldberg maximal operator, as we wanted to show. ∎
The analogous result for the operator $\widetilde{M}_{W}$ is a consequence of Theorem 1. Note here that the exponent that we get for $W]_{A_{p}}$ in this case is larger than that of Theorem 1. This is different from what is known in the scalar context, where the sharp exponent for the Christ–Goldberg maximal auxiliary operator is the same as for the Christ–Goldberg maximal operator (see [26, Lemma 3.5]). The reason for this larger exponent in our case is that first we bound $\widetilde{M}_{W}f$ by the Hardy–Littlewood maximal function $M(M_{W}f)$ of the Christ–Goldberg maximal operator, and then we apply Theorem 1. In the scalar setting, the proof of the analogous result relies on the reverse Hölder inequality for $A_{p}$ weights. We do not know if the exponent appearing in Theorem 6 can be improved.
@@
An application of the classical Fefferman–Stein vector valued inequalities for the maximal operator followed by the use of Theorem 1 yields the desired result. ∎
@@ 4 Two matrix weighted biparameter product BMO
In this second main part of the present work dyadic grids of cubes in $\mathbb{R}^{n}$, product dyadic grids in $\mathbb{R}^{n}\times\mathbb{R}^{m}$ and the associated Haar systems play a fundamental role. See Section 2 for the notation.
Fix a product dyadic grid $\bm{\mathcal{D}}=\mathcal{D}^{1}\times\mathcal{D}^{2}$ in $\mathbb{R}^{n}\times\mathbb{R}^{m}.$ Let $1<p<\infty,$ and let $U,V$ be biparameter $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic $A_{p}$ weights on $\mathbb{R}^{n}\times\mathbb{R}^{m}.$
Denote by $B$ any sequence $\{B_{R}^{\varepsilon}\}_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}$ in $\mathrm{M}_{d}(\mathbb{C}).$ We emphasize that $\mathcal{E}$ stands for the set of biparameter signatures, $\mathcal{E}=\mathcal{E}^{1}\times\mathcal{E}^{2},$ where $\mathcal{E}^{1}=\{0,1\}^{n}\setminus\{(1,\ldots,1)\}$ and $\mathcal{E}^{2}=\{0,1\}^{m}\setminus\{(1,\ldots,1)\}.$ We define
@@
$|\mathcal{V}_{R}P\mathcal{U}^{-1}_{R}|=|\mathcal{U}^{-1}_{R}P^{\ast}\mathcal{V}_{R}|\sim_{p,d}|\mathcal{U}^{\prime}_{R}P^{\ast}(\mathcal{V}^{\prime}_{R})^{-1}|,\qquad\forall R\in\bm{\mathcal{D}},\penalty 10000\ \forall P\in\mathrm{M}_{d}(\mathbb{C}).$
@@ 4.1 $\mathrm{H}^{1}$-BMO duality
The main goal of this subsection is to prove Theorem 2. We split the proof in Propositions 7 and 8 below, each treating one of the bounds in the equivalence stated in Theorem 2.
@@
This is the direct biparameter analog of the one-parameter two matrix weighted $\mathrm{H}^{1}$ norm defined in [28]. It is easy to check that $(\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p),\|\,\cdot\,\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)})$ is a Banach space.
@@
Let $B\in\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p).$ Then, the linear functional $\ell_{B}:\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)\rightarrow\mathbb{C}$ given by
@@
is well-defined and bounded with $\|\ell_{B}\|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\emph{BMO}_{\emph{prod},\bm{\mathcal{D}}}(U,V,p)}.$
@@
The proof of this result is essentially the same as in the one-parameter case for $p=2,$ which corresponds to the first half of the proof of [28, Theorem 1.3]. The necessary tools for the argument have been developed previously in other papers. We include the full proof for the reader’s convenience, indicating were to find the necessary facts to complete the arguments in this more general context.
Let $\Phi\in\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)$ be arbitrary. We show that the sum
$\ell_{B}(\Phi)\coloneq\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}\mathrm{tr}((B_{R}^{\varepsilon})^{\ast}\Phi_{R}^{\varepsilon})$
@@
$|\ell_{B}(\Phi)|\lesssim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|B\|_{\text{BMO}_{\text{prod},\bm{\mathcal{D}}}(U,V,p)}\|\Phi\|_{\mathrm{H}^{1}_{\bm{\mathcal{D}}}(U,V,p)}.$
@@
$\lesssim_{d}\sum_{\begin{subarray}{c}R\in\bm{\mathcal{D}}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|.$
The passage from the second line to the third one can be done using that $|\mathcal{V}^{-1}_{R}e|\leq|\mathcal{V}^{\prime}_{R}e|$ for any $e\in\mathbb{C}^{d}.$ In particular, if this fact is applied to the matrix norm $|\mathcal{V}_{R}^{-1}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|,$ one gets the stated inequality with an implicit constant depending on $d.$ Set now
@@
$\mathcal{B}_{k}\coloneq\left\{R\in\bm{\mathcal{D}}:\penalty 10000\ |R\cap\Omega_{k+1}|\leq\frac{1}{2}|R|<|R\cap\Omega_{k}|\right\},\qquad k\in\mathbb{Z},$
@@
Clearly $\Omega_{k}\subseteq\widetilde{\Omega}_{k}$ up to a set of zero measure, and in fact $|\widetilde{\Omega}_{k}|\sim|\Omega_{k}|,$ for all $k\in\mathbb{Z}.$ It is also obvious that $R\subseteq\widetilde{\Omega}_{k},$ for all $R\in\mathcal{B}_{k},$ for all $k\in\mathbb{Z}.$ Using the fact that $F\in\mathrm{L}^{1}(\mathbb{R}^{n+m}),$ it is easy to see that that for every $R\in\bm{\mathcal{D}}$ with $\Phi_{R}^{\varepsilon}\neq 0$ for some $\varepsilon\in\mathcal{E}$ there exists $k\in\mathbb{Z}$ such that $R\in\mathcal{B}_{k}.$ Thus, we have
$|\ell_{B}(\Phi)|$ $\lesssim_{p,d}\sum_{k\in\mathbb{Z}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|\mathcal{V}^{\prime}_{R}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|\cdot|\mathcal{V}_{R}B_{R}^{\varepsilon}\mathcal{U}_{R}^{-1}|$
@@
This will be enough to conclude the proof, because assuming it we will get that the sum over $k$ of the terms in the left-hand side of (4.1) will be bounded by
$V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\sum_{k\in\mathbb{Z}}2^{k}|\widetilde{\Omega}_{k}|\sim[V]_{A_{p},\bm{\mathcal{D}}}^{2/p}\|F\|_{\mathrm{L}^{1}(\mathbb{R}^{n+m})}.$
@@
by the definition of $\Omega_{k+1}.$ Moreover, denoting by $e_{1},\ldots,e_{d}$ the standard basis vectors in $\mathbb{C}^{d},$ we have
$\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}F(x)^{2}\,\mathrm{dm}(x)\geq\int_{\widetilde{\Omega}_{k}\setminus\Omega_{k+1}}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\frac{\mathbf{1}_{R}(x)}{|R|}\,\mathrm{dm}(x)$ (4.2)
$=\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}|^{2}\,\mathrm{dm}(x)$ (4.3)
$\sim_{d}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|^{2}\,\mathrm{dm}(x)$ (4.4)
$\geq\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\frac{|R\setminus\Omega_{k+1}|}{|R|}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$ (4.5)
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R\setminus\Omega_{k+1}|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$ (4.6)
$\sim\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R\setminus\Omega_{k+1}}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}.$ (4.7)
@@
$w]_{A_{p^{\prime}},\bm{\mathcal{D}}}\lesssim_{p,d}[V^{\prime}]_{A_{p^{\prime}},\bm{\mathcal{D}}}\sim_{p,d}[V]_{A_{p},\bm{\mathcal{D}}}^{p^{\prime}/p},$
because $V^{-1/p}$ is a $(d\times d)$ matrix $\bm{\mathcal{D}}$-dyadic biparameter $A_{p}$ weight (see for example Lemma 3.2 in [14]). It is then well-known that
$w^{1/p^{\prime}}]_{A_{2},\bm{\mathcal{D}}}\leq[w]_{A_{p^{\prime}},\bm{\mathcal{D}}}^{1/p^{\prime}}$
@@
(see for example Subsection 2.3.3 in [31]). Using Jensen’s inequality and the definition of the $A_{2}$ characteristic, it follows that
@@
Thus, the last part of (4.2) has the lower bound
$[V]_{A_{p},\bm{\mathcal{D}}}^{-2/p}\sum_{j=1}^{d}\sum_{\begin{subarray}{c}R\in\mathcal{B}_{k}\\ \varepsilon\in\mathcal{E}\end{subarray}}\left(\frac{1}{|R|}\int_{R}|V(x)^{-1/p}\Phi_{R}^{\varepsilon}\mathcal{U}_{R}e_{j}|\,\mathrm{dm}(x)\right)^{2}$